# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v2

@@ I-A Background of Large Model Based Agents
With the rise of large models (LMs) [5], including large language models (LLMs) such as OpenAI GPT-4, Google PaLM 2, and Microsoft Copilot, as well as large vision models (LVMs) such as DALL-E 2, LM-based agents (also known as LM agents or agentic LMs [6]) are unlocking new possibilities [7]. LM agents significantly enhance the inherent capabilities of AI systems, providing a versatile foundation for the next-generation AI agents [8]. Serving as the “brain” of AI agents, LMs empower them with advanced capabilities in human-machine interaction (HMI), few/zero-shot planning, contextual understanding, knowledge retention, and general-purpose task solving across physical, virtual, or mixed-reality environments [9, 10, 11]. LM agents generally fall into two categories:
  • Software-form LM agents, such as AutoGPT [12] and AutoGen [13], can autonomously interpret human instructions and use various tools (e.g., search engines and external APIs) to gather information and complete intricate tasks [14]. For instance, as shown in Fig. 2(a), an LM-powered virtual assistant can generate personalized travel plans, set reminders, and manage tasks while continuously learning and adapting in dynamic environments.
@@
LM agents are recognized as a significant step towards achieving artificial general intelligence (AGI) and have been widely applied across fields such as web search [16], recommendation systems [17], virtual assistants [18, 19], Metaverse gaming [20], robotics [15], autonomous vehicles [21], and automated attack penetration [22]. As reported by MarketsandMarkets [23], the worldwide market for autonomous AI and autonomous agents was valued at USD 4.8 billion in 2023 and is projected to grow at a CAGR of 43%, reaching USD 28.5 billion by 2028. LM agents have attracted global attention, and leading technology giants including Google, OpenAI, Microsoft, IBM, AWS, Oracle, NVIDIA, and Baidu are venturing into the LM agent industry.
Figure 2: Use cases of LM agents: (a) a software-form LM agent acting as the virtual assistant; (b) an embodied LM agent serving as the home cleaner. An LM agent, either in software or embodied form, generally consists of four key components: interaction, planning, action, and memory.
Figure 3: Overview of LM agent networks. Each LM agent includes two parts: (i) the brain located in the cyberspace powered by LMs such as GPT-4o, PaLM 2, and Copilot; and (ii) the physical or digital body such as autonomous vehicle, robot dog, UAV, and digital human. Within each LM agent, the cyber brain synchronizes with its physical/digital body via intra-agent communications. LM agents communicate with one another in the cloud to share task-oriented, knowledge, and inference results via inter-agent communications, establishing a network of interconnected intelligence. Each LM agent dynamically interact with other agents, virtual/real environments, and humans. The brain of each LM agent can be deployed either as a standalone entity or in a hierarchical manner across various platforms such as cloud servers, edge servers, or end devices.
@@ I-B Motivation of LM Agent Networks
Beyond the intelligence of single LM agents, connected LM agents can form LM agent networks to cooperatively address complex tasks beyond the capacity of individual agents [24]. As depicted in Fig. 3, within LM agent networks, connected LM agents can freely share sensory data, task-oriented knowledge, and LM inference results. By leveraging collective computational power and shared expertise from specialized LM agents, it fosters cooperative decision-making and collective intelligence. For instance, in autonomous driving, connected autonomous vehicles, acting as LM agents, share real-time sensory data, coordinate movements, and negotiate passage at intersections to optimize traffic flow and enhance safety. Each LM agent consists of an LM-powered brain (hosted in the cloud, edge, or end devices) and a physical or digital body. For embodied LM agents, a tiny local LM within the physical body handles local inference, while compute-intensive tasks are offloaded to the cloud LM, enabling collaborative computation between the cloud brain and the local brain. Each LM agent can interact with other agents, virtual/real environments, and humans. Communication within LM agent networks occurs via two primary modes: (i) intra-agent communications for seamless synchronization of status, data, and knowledge between the brain and body to enable real-time functionality; and (ii) inter-agent communications for efficient information exchange and computational coordination among multiple LM agents.
@@ I-C Motivation of Securing LM Agent Networks
1) Security vulnerabilities. LM agents are susceptible to “hallucinations”, where their LMs generate outputs that are plausible but incorrect and not grounded in reality, potentially spreading misinformation in multi-agent environments [34]. Hallucinations can compromise the reliability of decision-making, cause task failures, and pose risks to both physical entities and humans. Ensuring the integrity of intermediate inference results in collaborative systems is crucial, as biased or manipulated inputs can lead to unfair or erroneous outcomes [36]. Attacks such as adversarial manipulations [32], poisoning [37], and backdoors [38] exacerbate these vulnerabilities by allowing malicious actors to manipulate inputs and deceive the models. In collaborative environments, agent poisoning behaviors [33], where malicious agents disrupt the operation of others, can undermine the collaborative systems. Additionally, integrating LM agents into cyber-physical-social systems (CPSS) expands the attack surface, providing more opportunities for malicious exploitation within interconnected systems.
2) Privacy breaches. The extensive data collection processes and data memorization capabilities of LM agents raise significant risks related to data breaches and unauthorized access. These agents often process large volumes of personal and sensitive business information in both to-customer (ToC) and to-business (ToB) applications, which raises concerns regarding data storage, processing, and sharing [39]. Furthermore, LMs can inadvertently memorize sensitive details from training data and past interactions, potentially exposing private information during subsequent interactions [35]. These privacy risks are exacerbated in multi-agent collaborations, where LM agents may unintentionally leak sensitive information about users, other agents, or their internal operations during communication and task execution.
@@ I-D Related Surveys and Our Contributions
Recently, LM agents have garnered significant interest across academia and industry, leading to a variety of research exploring their potential from multiple perspectives. Notable survey papers in this field are as below. Andreas et al. [40] present a toy experiment for AI agent construction and case studies on modeling communicative intentions, beliefs, and desires. Wang et al. [41] identify key components of LLM-based autonomous agents (i.e., profile, memory, planning, and action) and the subjective and objective evaluation metrics. Besides, they discuss the applications of LLM agents in engineering, natural science, and social science. Xi et al. [10] present a general framework for LLM agents consisting of brain, action, and perception. Besides, they explore applications in single-agent, multi-agent, and human-agent collaborations, as well as agent societies. Zhao et al. [42] offer a systematic review of LLMs in terms of pre-training, adaptation tuning, utilization, and capacity assessment. Besides, background information, mainstream technologies, and critical applications of LLMs are introduced. Xu et al. [43] provide a tutorial on key concepts, architecture, and metrics of edge-cloud AI-generated content (AIGC) services in mobile networks, and identify several use cases and implementation challenges. Huang et al. [2] offer a taxonomy of AI agents in virtual/physical environments, discuss cognitive aspects of AI Agents, and survey the applications of AI agents in robotics, healthcare, and gaming. Cheng et al. [11] review key components of LLM agents (including planning, memory, action, environment, and rethinking) and their potential applications. Planning types, multi-role relationships, and communication methods in multi-agent systems are also reviewed. Masterman et al. [9] provide an overview of single-agent and multi-agent architectures in industrial projects and present the insights and limitations of existing research. Guo et al. [44] discuss the four components (i.e., interface, profiling, communication, and capabilities acquisition) of LLM-based multi-agent systems and present two lines of applications in terms of problem solving and world simulation. Durante et al. [45] introduce multimodal LM agents and a training framework including learning, action, congnition, memory, action, and perception. They also discuss the different roles of agents (e.g., embodied, simulation, and knowledge inference), as well as the potentials and experimental results in different applications including gaming, robotics, healthcare, multimodal tasks, and NLP. Hu et al. [20] outline six key components (i.e., perception, thinking, memory, learning, action, and role-playing) of LLM-based game agents and review existing LLM-based game agents in six types of games. Xu et al. [5] provide a comprehensive survey of enabling architectures and challenges for LM agents in gaming. Qu et al. [46] provide a comprehensive survey on integrating mobile edge intelligence (MEI) with LLMs, emphasizing key applications of deploying LLMs at the network edge along with state-of-the-art techniques in edge LLM caching, delivery, training, and inference.
@@
Existing surveys on LM agents primarily focus on the general framework design of LLM agents and multi-agent systems, as well as their potential in specific applications. Distinguished from the above-mentioned existing surveys, this survey focuses on the networking aspect of LM agents, including the general architecture, enabling technologies, key characteristics, and collaboration paradigms to construct networked systems of LM agents within physical, virtual, or mixed-reality environments. Additionally, with the advances of LM agents, it is urgent to examine their security and privacy in future AI agent systems. This work comprehensively reviews the security and privacy issues of LM agents and discusses both existing and potential defense mechanisms, which are overlooked in previous surveys. Table I compares the contributions of our survey with previous related surveys in the field of LM agents.
TABLE I: A Comparison of Our Survey with Relevant Surveys
@@
Review on background, pre-training, adaptation tuning,
evaluation & utilization, capacity assessment, and critical
applications of LLMs.
2024 [43]
@@
Discussions on taxonomy of AI agents in virtual/physical
environments, cognitive aspects, and applications in robots,
healthcare, and gaming.
2024 [11]
@@
2024 [9]
Overview of single-agent and multi-agent architectures in
industrial projects and insights and limitations of research.
@@
Survey the enabling architectures and key challenges of LM
agents for games.
2024 [46]
@@
In this paper, we present a systematic review of the state-of-the-arts in both single and connected LM agents, focusing on cooperation paradigms, security and privacy threats, existing and potential countermeasures, and future trends. Our survey aims to 1) provide a broader understanding of how LM agents work and how they cooperate in multi-agent scenarios, 2) examine the scope and impact of security and privacy challenges associated with LM agents and their interactions, and 3) highlight effective strategies and solutions for defending against these threats to safeguard LM agents in various intelligent applications. The main contributions of this work are four-fold.
@@
  • We comprehensively analyze existing and potential security and privacy threats, their underlying mechanisms, categorization, and challenges for both single and connected LM agents. We also review state-of-the-art countermeasures and examine their feasibility in securing LM agent networks.
@@ I-E Paper Organization
The remainder of this paper is organized as below. Section II gives an overview of single LM agents, while Section III presents the fundamentals for networking LM agents. Section IV and Section V introduce the taxonomy of security and privacy threats to LM agents, respectively, along with state-of-the-art countermeasures. Section VI outlines open research issues and future directions in the field of LM agents. Finally, conclusions are drawn in Section VII. Fig. 4 depicts the organization structure of this survey.
TABLE II: Progress of Standards for LM Agents
@@
The natural language interface is defined, including various protocols and guidelines that enable applications
and agents to effectively communicate with LLM agents, such as API syntax and semantics.
@@ II An Overview of Large Model Agents
In this section, we first introduce existing standards of LM agents. Then, we discuss the operating system (OS) and constructing modules of LM agents. Next, we introduce typical prototypes and discuss modern applications of LM agents.
Figure 5: Illustration of the OS architecture of LM agents [47, 48].
@@ II-A Standards of LM Agents
1) The IEEE SA - P3394 standard [^1] https://standards.ieee.org/ieee/3394/11377/, launched in 2023, defines natural language interfaces to facilitate communication between LLM applications, agents, and human users. This standard establishes a set of protocols and guidelines that enable applications and agents to effectively communicate with LLM agents. These protocols and guidelines include, but are not limited to, API syntax and semantics, voice and text formats, conversation flow, prompt engineering integration, LLM thought chain integration, as well as API endpoint configuration, authentication, and authorization for LLM plugins. The standard is expected to advance technological interoperability, promote AI industry development, enhance the practicality and efficiency of LMs, and improve AI agents’ functionality and user experience.
2) The IEEE SA - P3428 standard [^2] https://standards.ieee.org/ieee/3428/11489/, launched in 2023, aims to develop standards for LLM agents in educational applications. The primary goal is to ensure the interoperability of LLM agents across both open-source and proprietary systems. Key areas of focus include the integration of LLMs with existing educational systems and addressing technical and ethical challenges. This includes ensuring that LLMs can seamlessly interact with other AI components, such as adaptive instructional systems (AIS), while addressing issues of bias, transparency, and accountability within educational contexts. The standard is intended to support the widespread and effective application of LLMs in the field of education, thereby enabling more personalized, efficient, and ethically sound AI-driven educational experiences.

### v3

@@ I-A Background of Large Model Based Agents
With the rise of large models (LMs), including large language models (LLMs) such as OpenAI GPT-4, DeepSeek-R1, and Google PaLM 2, as well as multi-modal vision-language models (VLMs) such as Sora, LM-based agents (also known as LM agents or agentic LMs [5]) are unlocking new possibilities. LM agents significantly enhance the inherent capabilities of AI systems, providing a versatile foundation for the next-generation AI agents [6]. Serving as the “brain” of AI agents, LMs empower them with advanced capabilities in human-machine interaction (HMI), few/zero-shot planning, contextual understanding, knowledge retention, and general-purpose task solving across physical, virtual, or mixed-reality environments [2, 7]. LM agents generally fall into two categories:
  • Virtual LM agents, such as AutoGPT [8] and AutoGen [9], can autonomously interpret human instructions and use various tools (e.g., search engines and external APIs) to gather information and complete intricate tasks [10]. For instance, as shown in Fig. 2(a), an LM-powered personal assistant can generate personalized travel plans, set reminders, and manage tasks while continuously learning and adapting in dynamic environments.
@@
LM agents are recognized as a significant step towards achieving artificial general intelligence (AGI) and have been widely applied across fields such as web search [12], virtual assistants [13], Metaverse gaming [14], robotics [11], autonomous vehicles [15], and automated attack penetration [16]. As reported by MarketsandMarkets [17], the worldwide market for autonomous AI and autonomous agents was valued at USD 4.8 billion in 2023 and is projected to grow at a CAGR of 43%, reaching USD 28.5 billion by 2028. LM agents have attracted global attention, and leading technology giants including Google, OpenAI, Microsoft, IBM, AWS, Oracle, NVIDIA, and Baidu are venturing into the LM agent industry.
Figure 2: Use cases of LM agents: (a) a virtual LM agent acting as the personal assistant; (b) an embodied LM agent serving as the home cleaner. An LM agent, either in software or embodied form, generally consists of four key components: interaction, planning, action, and memory.
Figure 3: Overview of LM agent networks. Each LM agent consists of an LM-powered brain and a physical or digital body. Within each LM agent, the brain synchronizes with its body via intra-agent communications. LM agents communicate with one another in the cloud to share task-oriented, knowledge, and inference results via inter-agent communications, establishing a network of interconnected intelligence.
TABLE I: Summary of Key Abbreviations in Alphabetical Order
Abbr. Definition Abbr. Definition Abbr. Definition
ACL Agent Communications Language AI Artificial Intelligence AIGC AI-Generated Content
AIS Adaptive Instructional Systems AR/VR/MR Augmented/Virtual/Mixed Reality ASR Attack Success Rate
BKI Backdoor Keyword Identification CoT Chain-of-Thought CPSS Cyber-Physical-Social Systems
CroPA Cross-Prompt Attack DEPS Describe-explain-plan-select DHTs Distributed Hash Tables
DoS Denial-of-Service DP Differential Privacy FFN Feed-Forward Network
FPR False Positive Rate GoT Graph-of-Thought HMI Human-Machine Interaction
IP Intellectual Property KG Knowledge Graph LFU Least Frequently Used
LLMs Large Language Models LMs Large Models LRU Least Recently Used
LVMs Large Vision Models MAD Multi-Agent Debate MARL Multi-Agent Reinforcement Learning
MEI Mobile Edge Intelligence MFA Multi-Factor Authentication MIAs Membership Inference Attacks
MITM Man-In-The-Middle MLMs Masked Language Models MoE Mixture-of-Experts
MSE Mean Squared Error NDN Named Data Networking NLP Natural Language Processing
NMT Neural Machine Translation NSFW Not-Safe-For-Work OS Operating System
PEFT Parameter-Efficient Fine-Tuning PII Personally Identifiable Information PoT Plan-of-Thought
PPI Personal Preference Information QoS Quality-of-Service RAG Retrieval-Augmented Generation
ROS Robot Operating System SKGR Self-Knowledge-Guided Retrieval SL Split Learning
SPP Solo Performance Prompting TMM Transferable Multi-Modal ToB To-Business
ToC To-Customer ToT Tree-of-Thought TPSP Third-Party Service Provider
SHAP Shapley Additive Explanations VLMs Vision-Language Models XAI Explainable Artificial Intelligence
@@ I-B Motivation of LM Agent Networks
Beyond the intelligence of single LM agents, connected LM agents can form LM agent networks to cooperatively address complex tasks beyond the capacity of individual agents [18]. As depicted in Fig. 3, within LM agent networks, connected LM agents can freely share sensory data, task-oriented knowledge, and LM inference results. By leveraging collective computational power and shared expertise from specialized LM agents, it fosters cooperative decision-making and collective intelligence. For instance, in autonomous driving, connected autonomous vehicles, acting as LM agents, share real-time sensory data, coordinate movements, and negotiate passage at intersections to optimize traffic flow and enhance safety. Each LM agent includes two parts: (i) the brain located in the cyberspace powered by LMs such as GPT-4o, PaLM 2, and DeepSeek-R1; and (ii) the physical or digital body such as autonomous vehicle, robot dog, UAV, and digital human. For embodied LM agents, a tiny local LM within the physical body handles local inference, while compute-intensive tasks are offloaded to the cloud LM, enabling collaborative computation between the cloud brain and the local brain. Each LM agent can interact with other agents, virtual/real environments, and humans. The brain of each LM agent can be deployed either as a standalone entity or in a hierarchical manner across various platforms such as cloud servers, edge servers, or end devices. Communication within LM agent networks occurs via two primary modes: (i) intra-agent communications for seamless synchronization of status, data, and knowledge between the cyber brain and physical/digital body to enable real-time functionality; and (ii) inter-agent communications for efficient information exchange and computational coordination among multiple LM agents.
@@ I-C Motivation of Securing LM Agent Networks
1) Security/reliability vulnerabilities. LM agents are susceptible to “hallucinations”, where their LMs generate outputs that are plausible but incorrect and not grounded in reality, potentially spreading misinformation in multi-agent environments [28]. Hallucinations can undermine the reliability of decision-making, cause task failures, and pose risks to both physical systems and human safety. Ensuring the integrity of intermediate inference results in collaborative systems is crucial, as biased or manipulated inputs can produce unfair or erroneous outcomes [30]. Additionally, attacks such as adversarial manipulations [26], poisoning [31], and backdoors [32] exacerbate these vulnerabilities by allowing malicious actors to manipulate inputs and deceive models. In collaborative settings, agent poisoning behaviors can undermine the collaborative systems [27], as malicious agents can disrupt the operation of others. Furthermore, integrating LM agents into cyber-physical-social systems (CPSS) expands the attack surface, increasing opportunities for malicious exploitation within interconnected systems.
2) Privacy breaches. The extensive data collection processes and data memorization capabilities of LM agents pose significant privacy risks related to data misuse and unauthorized access. These agents often handle large volumes of personal and business-sensitive data in both to-customer (ToC) and to-business (ToB) applications, raising concerns about secure storage, processing, and sharing [33]. Furthermore, LMs can inadvertently memorize sensitive information from training data or past interactions, potentially exposing them in subsequent interactions [29]. These privacy risks are exacerbated in multi-agent collaborations, where LM agents may unintentionally leak confidential user data, internal operational details, or proprietary business information during communication and task execution.
@@ I-D Related Surveys and Our Contributions
Recently, LM agents have garnered significant interest across academia and industry, leading to a variety of research exploring their potential from multiple perspectives. Notable survey papers in this field are as below. Andreas et al. [34] present a toy experiment for AI agent construction and case studies on modeling communicative intentions, beliefs, and desires. Wang et al. [35] identify key components of LLM-based autonomous agents (i.e., profile, memory, planning, and action) and the subjective and objective evaluation metrics. Besides, they discuss the applications of LLM agents in engineering, natural science, and social science. Xi et al. [2] present a general framework for LLM agents consisting of brain, action, and perception. Besides, they explore applications in single-agent, multi-agent, and human-agent collaborations, as well as agent societies. Xu et al. [36] provide a tutorial on key concepts, architecture, and metrics of edge-cloud AI-generated content (AIGC) services in mobile networks, and identify several use cases and implementation challenges.
Cheng et al. [7] review key components of LLM agents (including planning, memory, action, environment, and rethinking) and their potential applications. Planning types, multi-role relationships, and communication methods in multi-agent systems are also reviewed. Guo et al. [37] discuss the four components (i.e., interface, profiling, communication, and capabilities acquisition) of LLM-based multi-agent systems and present two lines of applications in terms of problem solving and world simulation. Durante et al. [38] introduce multimodal LM agents and a training framework including learning, action, congnition, memory, action, and perception. They also discuss the different roles of agents (e.g., embodied, simulation, and knowledge inference), as well as the potentials and experimental results in different applications including gaming, robotics, healthcare, multimodal tasks, and NLP. Hu et al. [14] outline six key components (i.e., perception, thinking, memory, learning, action, and role-playing) of LLM-based game agents and review existing LLM-based game agents in six types of games. Qu et al. [39] provide a comprehensive survey on integrating mobile edge intelligence (MEI) with LLMs, emphasizing key applications of deploying LLMs at the network edge along with state-of-the-art techniques in edge LLM caching, delivery, training, and inference.
For the security and privacy of LLMs and agents, Yao et al. [40] provide a comprehensive review of LLMs’ dual role in cybersecurity, examining their potential to enhance security and privacy (“the good”), the risks and threats they introduce (“the bad”), and their inherent vulnerabilities (“the ugly”). Friha et al. [41] comparatively review recent optimization and autonomy methods for resource-constrained edge environments, while reviewing security concerns, trust issues, and mitigation strategies for LLM-powered MEI. Das et al. [42] explore the security and privacy challenges of LLMs, analyzing application-specific risks across various domains, including transportation, education, and healthcare. They also evaluate the scope of LLM vulnerabilities, emerging attack vectors, and potential defense strategies.
@@
Existing surveys on LM agents primarily focus on the general framework design of LLM agents and multi-agent systems, as well as their potential in specific applications. Distinguished from the above-mentioned existing surveys, this survey focuses on the networking aspect of LM agents, including the general architecture, enabling technologies, key characteristics, and collaboration paradigms to construct networked systems of LM agents within physical, virtual, or mixed-reality environments. Additionally, with the advances of LM agents, it is urgent to examine their security and privacy in future AI agent systems. This work comprehensively reviews the security and privacy issues of LM agents and discusses both existing and potential defense mechanisms, which are overlooked in previous surveys. Table II compares the contributions of our survey with previous related surveys in the field of LM agents.
TABLE II: A Comparison of Our Survey with Relevant Surveys
@@
2024 [40]
Survey on LLMs’ potential to enhance security & privacy,
risks they introduce, and their inherent vulnerabilities.
2024 [41]
Survey on optimization and autonomy techniques for edge
environments, review security concerns, trust issues, and
mitigation strategies for LLM-powered MEI applications.
2025 [42]
Explore security and privacy challenges of LLMs, analyze
application-specific LLM risks, and evaluate the scope of
LLM vulnerabilities, emerging attack vectors, and potential
defense strategies.
@@
In this paper, we present a systematic review of the state-of-the-arts in both single and connected LM agents, focusing on cooperation paradigms, security and privacy threats, existing and potential countermeasures, and future trends. Our survey aims to 1) provide a broader understanding of how LM agents work and how they cooperate in multi-agent scenarios, 2) examine the scope and impact of security, privacy, and reliability challenges associated with LM agents and their interactions, and 3) highlight effective strategies and solutions for defending against these threats to safeguard LM agents in various intelligent applications. The main contributions of this work are four-fold.
@@
  • We comprehensively analyze existing and potential security, privacy, and reliability threats, their underlying mechanisms, categorization, and challenges for both single and connected LM agents. We also review state-of-the-art countermeasures and examine their feasibility in securing LM agent networks.
@@ I-E Paper Organization
The remainder of this paper is organized as below. Section II gives an overview of single LM agents, while Section III presents the fundamentals for networking LM agents. Section IV and Section V introduce the taxonomy of security and privacy threats to LM agents, respectively, along with state-of-the-art countermeasures. Section VII outlines open research issues and future directions in the field of LM agents. Finally, conclusions are drawn in Section VIII. Fig. 4 depicts the organization structure of this survey, and Table I summarizes the key acronyms.
TABLE III: Progress of Standards for LM Agents
@@
The natural language interface is defined, including various protocols and guidelines to facilitate seamless
and efficient interaction between APPs, agents, and LLM-powered systems.
@@ II An Overview of Large Model Agents
In this section, we first introduce existing standards of LM agents. Then, we discuss the operating system (OS) and constructing modules of LM agents. Next, we introduce typical prototypes and discuss modern applications of LM agents. Fig. 5 shows the organization structure of this section.
Figure 5: Organization structure of the Section II.
@@ II-A Standards of LM Agents
1) The IEEE SA - P3394 standard [^1] https://standards.ieee.org/ieee/3394/11377/, launched in 2023, defines natural language interfaces to enhance communication between LLM applications, agents, and human users. It establishes a series of protocols and guidelines to facilitate seamless and efficient interaction between applications, agents, and LLM-powered systems. These protocols and guidelines include, but are not limited to, API syntax and semantics, voice and text formats, conversation flow, prompt engineering integration, LLM thought chain integration, as well as API endpoint configuration, authentication, and authorization for LLM plugins. The standard is expected to advance technological interoperability, promote AI industry development, enhance the practicality and efficiency of LMs, and improve AI agents’ functionality and user experience.
2) The IEEE SA - P3428 standard [^2] https://standards.ieee.org/ieee/3428/11489/, launched in 2023, aims to develop standards for LLM agents in educational applications. The primary goal is to ensure the interoperability of LLM agents across both open-source and proprietary systems. Key areas of focus include the integration of LLMs with existing educational systems and addressing technical and ethical challenges. This includes ensuring that LLMs can seamlessly interact with other AI components, such as adaptive instructional systems (AIS), while addressing issues of bias, transparency, and accountability within educational contexts. The standard is intended to support the effective LLMs-empowered educational applications, thereby enabling more personalized, efficient, and ethically sound AI-driven educational experiences.
Figure 6: Illustration of the OS architecture of LM agents [43].